OpenStudy (anonymous):

Simplify. Please show the steps to arrive at your answer. 3^3/9*2-12+6

OpenStudy (mathteacher1729):

Which is the expression you are trying to simplify? \[\dfrac{3^3}{9}\cdot2-12+6\] or \[\dfrac{3^3}{9\cdot2-12}+6\] or \[\dfrac{3^3}{9\cdot2-12+6}\]

OpenStudy (anonymous):

the first one

OpenStudy (mathteacher1729):

Do you know "PEMDAS"?

OpenStudy (anonymous):

no